### v2

# UnMarker: A Universal Attack on Defensive Image Watermarking

Abstract
Reports regarding the misuse of Generative AI (GenAI) to create deepfakes are frequent. Defensive watermarking enables GenAI providers to hide fingerprints in their images and use them later for deepfake detection. Yet, its potential has not been fully explored. We present UnMarker— the first practical universal attack on defensive watermarking. Unlike existing attacks, UnMarker requires no detector feedback, no unrealistic knowledge of the watermarking scheme or similar models, and no advanced denoising pipelines that may not be available. Instead, being the product of an in-depth analysis of the watermarking paradigm revealing that robust schemes must construct their watermarks in the spectral amplitudes, UnMarker employs two novel adversarial optimizations to disrupt the spectra of watermarked images, erasing the watermarks. Evaluations against SOTA schemes prove UnMarker’s effectiveness. It not only defeats traditional schemes while retaining superior quality compared to existing attacks but also breaks semantic watermarks that alter an image’s structure, reducing the best detection rate to $43\%$ and rendering them useless. To our knowledge, UnMarker is the first practical attack on semantic watermarks, which have been deemed the future of defensive watermarking. Our findings show that defensive watermarking is not a viable defense against deepfakes, and we urge the community to explore alternatives.
1 Introduction
GenAI has amassed popularity due to its growing sophistication and accessibility [1, 2, 3]. This technology has solved problems in various fields [4, 5, 6, 7]. Yet, it comes with the risk of malicious entities using it to generate harmful images (deepfakes) [8, 9], proven by many reports of those being used in political smear campaigns [10] and non-consensual pornography [11].
Defenses first emerged as passive detectors [12, 13, 14, 15] that tell apart fake from real images, but these were proven ineffective, failing to generalize to unseen synthesis algorithms [16]. As training advanced generators is costly and prohibitive [17] in the absence of the extensive resources and high-quality data that only a few GenAI providers possess, adversaries typically resort to these providers’ APIs to generate harmful content. This recently led to the adoption of defensive watermarking as a countermeasure, which enables providers to embed hidden signatures in their images that can later be identified using designated detectors. Defensive watermarking has gained traction in the research community [18, 19, 20, 21, 22, 23, 24, 25] and the industry, as providers have already implemented and are testing such solutions (e.g., Google’s SynthID [26]) or pledged to do so soon [27] in commitments to the White House [28].
Defensive watermarking protects against limited adversaries, following the assumption that attackers lack the resources to train complex generators. Yet, existing schemes fail to demonstrate strong guarantees for this resistance beyond empirical robustness to naive manipulations. We devise UnMarker— the first universal (effective against all schemes), black-box (without access to the scheme’s parameters or similar systems), data-free (no additional data needed) and query-free (without detector feedback) attack.
Our motivating insight is that, for any scheme, there must be a universal carrier that embeds the watermark in all images. In layman’s terms, this carrier is the set of all attributes the detector may examine when verifying the watermark. This set must exist and is finite (shared among all images) as detectors are computationally limited. By analyzing the desired resistance to basic manipulations and the requirement that watermarks must naturally integrate with various contents, we deduce that this carrier can only be in spectral amplitudes. That is, a “conventionally” robust scheme must construct its watermarks by manipulating the frequencies at which the different pixels vary, forcing the collective distribution of these value shifts at different rates to follow a unique pattern. We can roughly even isolate the range in which watermarks are encoded: the more robust semantic schemes [18, 24], which alter the structure, target low frequencies while content-preserving non-semantic watermarks [21, 19, 23, 20] populate higher bands.
UnMarker disrupts these spectral amplitudes. To remove non-semantic watermarks, it directly modifies the image to optimize a designated loss, affecting the rates at which individual pixels vary w.r.t. their neighbors and impacting the distribution of high-frequency amplitudes. For semantic watermarks, we devise novel optimizable filters and a procedure to learn their weights to systemically alter the consistency at different regions, causing discrepancies at lower frequencies. Unlike known filters that rely on geometric heuristics, limiting them to minor distortions, UnMarker is guided by a perceptual loss to maximize the disturbance in visually non-critical areas without compromising quality. As a result, the carrier set is shifted away from the watermark’s distribution. By targeting the carrier, UnMarker abandons the dependence on a specific model, leaving it successful without knowledge of the scheme or detector feedback. Previous adversarial attacks [29, 30, 31, 32, 33, 34] all face this limitation, failing under realistic conditions.
We experiment with seven SOTA schemes from top venues. By combining the above techniques with mild cropping, UnMarker breaks them all, despite its black-box and query-free nature. We compare UnMarker with the recent VAEAttack [35] and DiffusionAttack [35, 29, 31], demonstrating far superior results, defeating semantic watermarks against which they fail and disputing misconceptions regarding these schemes’ robustness to limited adversaries [35, 31, 29]. Our results urge vendors to rethink their response to deepfakes. We make the following contributions:
  • We thoroughly analyze defensive watermarking and reveal trade-offs that make it an invalid countermeasure.
  • We employ novel spectral optimizations to devise UnMarker— the first universal data-free, black-box, and query-free attack against defensive watermarking [^1] https://github.com/andrekassis/ai-watermark.
  • We demonstrate UnMarker’s efficacy against seven SOTA schemes under restrictive settings, rendering them useless.
  • We compare UnMarker to the strongest query-free black-box attacks, attaining superior performance and quality.
  • Using a novel class of optimizable filters to incur structural disturbances, UnMarker is the first practical attack to defeat semantic watermarks, whose robustness has marked them as the ultimate candidates for trustworthy deepfake detection, showing this assumption is highly questionable.
2 Background & Related Work
2.1 Deepfake Detection
Passive Detectors. These systems [12, 13, 14, 15] identify deepfakes based on semantic incoherences [36, 37] and low-level artifacts [12, 14, 38, 39]. Advanced generators can produce high-quality outputs without such telltales, and manipulations can fool these detectors [40, 41, 42]. They also fail to generalize to unseen models [16].
Defensive Watermarking. It assumes attackers lack the resources to train high-quality generators. A watermarking scheme consists of watermark ($w$), encoder ($\mathcal{E}$) and detector ($\mathcal{D}$). In a general-purpose scheme [19, 24], $\mathcal{E}$ accepts any image $x$ and outputs a similar watermarked $x_{w}\!\!=\!\!\mathcal{E}(x,w)$. Given image $x^{\prime}$, $\mathcal{D}$ extracts a sequence $w^{\prime}\!\!=\!\!\mathcal{D}(x^{\prime})$ that will be almost identical to $w$ $iff$ $x^{\prime}$ was previously watermarked. Generator-specific approaches “watermark” generative models: $\mathcal{E}$ can be the generator itself trained to only produce watermarked images [21, 22] or an algorithm that modifies its weights [23, 20] or latent codes [18] s.t. images it produces embed the watermark. Heuristics-based watermarks [43, 44, 45] directly manipulate spectral representations, forming a signature in select frequencies. These are removable via post-processing [30, 20]. We focus on learning-based methods [19, 24, 21, 22, 23, 20, 18, 25], where $\mathcal{E}$ and $\mathcal{D}$ are DNNs as they resist post-processing.
Semantic vs. Non-Semantic: Non-semantic schemes take an image and produce a visually identical watermarked version. Aside from StegaStamp [24], general-purpose schemes are non-semantic. Generator-specific schemes, except TRW [18], are also non-semantic, as their training restricts them to output images identical to non-watermarked references. Semantic watermarks can influence the content. This term was popularized by Wen et al. with their introduction of TRW [18] that alters the latent codes used by text-to-image Latent Diffusion Models (LDMs) [46]. When providing a random code with a textual phrase, the LDM uses this code as a seed to produce an image matching the text (e.g, “a dog on a bike” results in an image of a dog riding a bike but these objects will change based on the seed). TRW injects a pattern in the Fourier transforms of all latent codes before generation. As a result, watermarked images will differ from non-watermarked ones but still satisfy the text. This leaves watermarks invisible but highly robust. StegaStamp [24] precedes TRW and operates similar to its general-purpose counterparts. Yet, it is less stealthy, causing structural changes (see Fig 2). Still, its strategically distributed distortion ensures minimal visibility. In defensive watermarking, users cannot observe non-watermarked images, further hiding these nuances and making StegaStamp highly applicable. While semantic watermarks primarily refer to TRW, previous work recognizes the similarities in both schemes, distinguishing them as more robust [35, 29, 31]. Thus, we refer to TRW and StegaStamp as semantic.
2.2 Attacks Against Defensive Watermarking
Regeneration Attacks. Proposed by Zhao et al. [35], they rely on Diffusion Models (DMs) [47] or Variational Autoencoders (VAEs) [48] (henceforth, the DiffusionAttack and VAEAttack) to introduce noise and then reconstruct images after the watermark has been eliminated. Saberi et al. [29] and WAVES [31] revisit these attacks. All these works verify non-semantic watermarks’ vulnerability, against which regeneration attacks offer theoretical guarantees. Yet, semantic schemes remain robust. While TRW experiences some degradation [31], it occurs at a considerable distortion where the attack is invalid (see §3). Regeneration is the only practical (black-box and query-free) attack but fails against semantic watermarks. It also relies on the denoising pipelines’ fidelity, often impacting quality (see §6.2.3). UnMarker defeats all watermarks while also abandoning the dependence on regeneration models, ensuring quality.
Adversarial Attacks. Realistically, attackers cannot access the scheme (see §3), making traditional adversarial attacks that target the decision boundaries of specific models ineffective. Thus, existing works make relaxing assumptions, failing to retain success under practical conditions.
WEVADE [30] attacks watermarked images by querying the detector and updating them accordingly until the watermark is not detected. Yet, detector feedback may not be available (see §3), making it ineffective. Lukas et al. [32] propose query-free attacks but assume unrealistic access to a generator with similar architecture and data (see §3). Saberi et al. [29] mount an attack on a surrogate detector, which transfers to the target. Training the substitute requires non-watermarked samples from the provider’s distribution, violating the defensive watermarking principles [31, 33] (see §3) as providers only release watermarked images and possess exclusive datasets. Still, it fails against the semantic StegaStamp without affecting quality. RD-IWAN [34] also requires access to non-watermarked provider images. An et al. [31] evaluate surrogate attacks specifically on semantic schemes, finding StegaStamp robust, while TRW is affected under unrealistic access to its autoencoder or images generated by it with different watermarks, which is impractical as any provider’s published image bear its single identifying watermark. They also prove Saberi et al.’s [29] surrogate attack unsuccessful absent non-watermarked provider images. Hu et al.’s [33] ensemble surrogate attack requires no non-watermarked provider images. Yet, the evaluations exclude semantic watermarks and only consider surrogates and targets of the same scheme, leading to sufficient commonalities for transferability and violating the black-box setting. It also requires extensive resources ($40\!\!-\!\!100$ surrogates), data, and large perturbations, making it impractical. UnMarker waives impractical assumptions, requires no data or prohibitive resources, and preserves quality.
3 Threat Model
Our framework considers two parties: Providers who develop and monetize these platforms, and Attackers abusing their access to generate malicious content.
Providers’ Goals. GenAI providers must ensure their images are not used to spread disinformation. Thus, they watermark generated images. For a suspicious image, they invoke the detector to test for the watermark, proving it fake.
Providers’ Capabilities. They have resources and data to train SOTA generators. As is common practice [1, 3], their APIs are restrictive, accepting queries (e.g., text) and returning watermarked images, hiding intermediate results.
Attacker’s Goals. Attackers wish to generate real-looking deepfakes by invoking the providers’ APIs and then using methods to 1) remove watermarks while 2) retaining quality.
Attacker’s Capabilities: Resources. Following the defensive watermarking assumptions (see §2), attackers lack high-quality data or extensive resources. Thus, training advanced models or surrogates for the providers’ generators [32] or detectors [29] is not possible. As some services (e.g., AWS) allow renting advanced hardware, attackers may access powerful machines. Yet, this can only be short-term due to high costs, not for training large models over long GPU hours.
Black-Box. As many providers keep their algorithms and datasets secret [49], attackers have no whitebox access to (similar) models, data, or any information about the system.
Query-Free. When providers restrict access to their detector to trusted corporations or employees, as confirmed by recent announcements [50], attacks relying on detector feedback [30] lose their efficacy. Given these announcements and defensive watermarking’s vulnerability when query access to the detector is publicly available (see §2.2), we follow previous work [33, 35, 29] and investigate this paradigm against query-free attacks where attackers cannot interact with the detector directly to observe its decisions.
4 Design Philosophy
Since the detector $\mathcal{D}$ must be able to verify the watermark in any image, it should know where to find this watermark, meaning there must be a set of attributes, namely the universal carrier, that can be measured in all images to make this decision. Our analysis will reveal that identifying this set is possible since the requirements from a robust scheme eliminate all options except the spectral amplitudes. Thus, any robust scheme will inject its watermarks in the collective magnitudes of the different frequency bands. In §5, we show how identifying this carrier enables attackers to mount destructive attacks that remove the watermark with minimal visual impact and no knowledge of the system.
4.1 Uncovering The Universal Carrier
Any computationally limited detector $\mathcal{D}$ can only monitor a finite set of attributes $\mathcal{\boldsymbol{P}}=\{p_{b}|\;b\in[\![B]\!]\}$ in all images. Deciding if an image is watermarked relies on whether (a subset of) $\mathcal{\boldsymbol{P}}$’s attributes follow a watermark property distribution $\mathcal{\boldsymbol{P}}_{w}$ that $\mathcal{D}$ associates with the watermark $w$. While $\mathcal{\boldsymbol{P}}$ is a set of attributes, $\mathcal{\boldsymbol{P}}_{w}$ is a distribution over $\mathcal{\boldsymbol{P}}$’s values, that identifies $w$. There may be several patterns $\mathcal{\boldsymbol{P}}_{w}$ pertaining to subsets of $\mathcal{\boldsymbol{P}}$. We term $\mathcal{\boldsymbol{P}}$ the universal carrier as it contains all the properties $w$ can influence. Although $\mathcal{\boldsymbol{P}}_{w}$ should allow watermarking different contents, its variance must still be sufficiently low to prevent detecting $w$ in non-watermarked images, restricting it to a reasonably narrow vicinity. Thus, introducing small deviations into $\mathcal{\boldsymbol{P}}$’s values can cause any watermarked image to violate this distribution $\mathcal{\boldsymbol{P}}_{w}$, removing the watermark without degrading quality. This destructive strategy requires no knowledge of $\mathcal{\boldsymbol{P}}_{w}$, making it feasible without feedback from $\mathcal{D}$ if the carrier $\mathcal{\boldsymbol{P}}$ is correctly identified. To identify $\mathcal{\boldsymbol{P}}$, we analyze widely-accepted requirements [18, 23, 19, 24, 21, 22, 20] from a robust scheme:
Robust Watermarking Requirements. Universal Applicabil- ity: The scheme must watermark images of various concepts and scenes, and the watermark must always be detectable.
Universal Stealthiness. The process cannot introduce unnatural artifacts that degrade quality. We do not require the watermark not to leave visible traces since semantic watermarks violate this concept by definition, influencing the structure. Yet, as users only observe watermarked images, quality becomes the only relevant metric for stealthiness.
Figure 1: Spectral analysis of two images. Both depict worldly objects of high consistency. Thus, the collective spectral magnitudes are distributed similarly as the low frequencies (center) corresponding to gradual variations are always far more dominant. Phases determine the spatial arrangement of the pixel value shifts that constitute these magnitudes to shape the content, making them extremely different.
Robustness to Manipulations. The watermark should resist known manipulations (e.g, cropping, blurring, compression).
4.1.1 Candidates
An image is characterized by its 1) spatial structure (pixel values), and 2) frequency distribution— the rates at which values alternate between neighboring pixels. Thus, the carrier must reside in these channels. Below, we present hypotheses, accompanied by an in-depth analysis, regarding the possibility of injecting watermarks in each of these channels, identifying a unique carrier capable of satisfying all demands. Our claims are corroborated by UnMarker’s empirical success in §6.
H1. Watermarks cannot be embedded in the spatial structure (pixel values). This is because they make for an unreliable carrier that experiences sporadic variations (in the absence of a watermark) between images based on their contents (different images assign extremely different values to their pixels). However, the scheme must always bring the carrier’s values to the same distribution $\mathcal{\boldsymbol{P}}_{w}$ for any watermarked image. When these values naturally (i.e., without the watermark) exhibit a high degree of content dependency, this is impossible without restricting the scheme to processing a small set of similar contents or leaving unnatural cues in many outputs corresponding to contents that would otherwise require these values in images portraying them to lie far away from the desired $\mathcal{\boldsymbol{P}}_{w}$. As this distance increases, the magnitudes of the changes needed to bring them to $\mathcal{\boldsymbol{P}}_{w}$ grow, failing to integrate naturally with the content. Restrictions violate applicability, while unnatural artifacts are not stealthy. The scheme may strategically place pixels encoding the watermark to accommodate large changes with minimal visibility. Two possibilities exist: either they are far apart to remain unnoticeable despite significantly differing from their environments or located in regions the content is unlikely to occupy (i.e., the background). Both are volatile and cannot withstand manipulations (e.g., cropping or noise), violating the robustness requirement (see §6.2.1).
Figure 2: Images watermarked by StableSignaturenon-semantic (top), and StegaStampsemantic (bottom). The rightmost figures display the differences between the original and watermarked images that correspond to the changes encoding the watermarks. StableSignature’s modifications are restricted to existing (high-frequency) edges such as wrinkles, hair, mustache, and intersections of multiple components. StegaStamp’s watermark is distributed across the image. The magnified area shows how it manipulates the consistency (texture), injecting gradual (low-frequency) changes that manifest as wrinkles at this location.
We now turn our focus to the frequency distribution. There are two options for influencing frequencies to inject watermarks: Locally or globally. We claim the following:
H2. Local changes to the pixel shift rates cannot reliably encode watermarks. The reason is that local modifications affect the rates at which pixels vary in select spatial regions. These share the above pitfall of strong dependency on the content, leaving only global spectral changes. The spectrum associates frequencies with two properties: magnitude and phase. Magnitudes represent the prevalence of shifts that occur at the specific rate across the image. Mathematically, any image is a superposition of periodic signals with varying frequencies occurring at different phases that determine their orientations, forming the image’s structure based on how they intersect. Accordingly, and similar to the spatial structure, we deduce that: H3. Phases are an unreliable carrier due to their high sensitivity to the content. This is the result of the phases varying drastically between images as the primary factor affecting the structure (see Fig. 1).
By process of elimination, hypotheses H1-H3 exclude all other candidates, leaving only one potential carrier: the global spectral magnitudes. These are collective measurements with little dependency on specific pixels, making them resistant to reasonable manipulations that would preserve the majority of energy (amplitudes) distribution. They are also distributed similarly for any realistic content: lower frequencies contain the majority of energy as they correspond to gradual changes that capture the inherent consistency (continuity) of the worldly objects images depict. Higher frequencies (rapid transitions) occur at edges (intersections of objects), making them far less prevalent (see Fig. 1). Thus, while amplitudes vary between images, they always follow a similar distribution, decaying as the frequencies grow. This stability other alternatives lack allows injecting watermarks via changes of similar magnitudes for any content, enabling them to be minimized s.t. they always integrate naturally.
4.1.2 Spectral Watermarks
The above entails that robust watermarks reside in the global spectral amplitudes. Based on the different schemes’ characteristics, we now present our main hypothesis, which serves as UnMarker’s foundation:
H4. Non-semantic schemes primarily inject watermarks in the global spectral amplitudes associated with high frequencies, while semantic watermarks mainly target low frequencies. Below, we analyze this claim:
Non-Semantic: As these schemes cannot change the content, these are typically injected in high frequencies since sufficiently manipulating lower bands s.t. the watermark is detected requires substantial changes due to the vast energy concentration in this range. To alter the collective distribution of high frequencies, polarization at existing edges can be controlled by adjusting the intensities of the various intersecting components to amplify or attenuate the transitions to their environments. This significantly impacts higher frequency amplitudes via small (invisible) modifications only due to the naturally low energy distribution in this range and by exploiting the characteristics of these regions (edges) that already exhibit high-frequency shifts to mask the watermark’s footprint. Fig. 2 (top) depicts this behavior for the non-semantic StableSignature [20]. We verified that other non-semantic schemes operate similarly.
Semantic: Since these schemes can influence the content, they will populate the consistency-defining low frequencies, making them more robust. As spectral amplitudes represent the collective contribution to pixel variations occurring at the relevant rates from all spatial locations, these schemes construct their outputs s.t. the needed changes are spread among different regions, reducing visual cues. Unlike non-semantic watermarks, they are not restricted to edges, affecting various objects’ structures and textures. Since the changes required to alter low-frequency magnitudes are large, their influence may be visible. Fig. 2 (bottom) portrays this for StegaStamp [24]. TRW’s [18] images may display different content from the original (non-watermarked), making illustrations uninformative (see §2.1). Yet, our results in §6.2.2 prove its watermarks are in low frequencies.
5 UnMarker
Now that we have identified the carrier, we present our universal attack. It follows from §4.1.1 (H4) that the carrier must be a set $\mathcal{\boldsymbol{F}}_{c}=\{|F_{b}|:\;b\in[\![B]\!]\}$ of spectral amplitudes and the watermark property distribution, denoted as ${\mathcal{\boldsymbol{F}}}_{w}$, is over these amplitudes. Let $\ell_{d}$ be a perceptual distance metric s.t. if $\ell_{d}(x,\;y)\leq t_{\ell_{d}}$ for some threshold $t_{\ell_{d}}$, then $x$ and $y$ are visually similar. We can now formalize watermark removal:
Watermark Removal.
(WR): Given watermarking algorithm $\mathcal{A}$ whose carrier frequency set is $\mathcal{\boldsymbol{F}}_{c}=\{|F_{b}|:\;b\in[\![B]\!]\}$ with watermark property distribution ${\mathcal{\boldsymbol{F}}}_{w}$, the watermark $w$ is removable from $x_{w}$ if there is an image $x_{nw}$ s.t. $\ell_{d}(x_{w},\;x_{nw})\leq t_{\ell_{d}}\land\mathcal{\boldsymbol{F}}_{c}(x_{nw})\not\sim\mathcal{\boldsymbol{F}}_{w}$.
Removing the watermark from $x_{w}$ amounts to finding a visually similar $x_{nw}$ for which $\mathcal{\boldsymbol{F}}_{c}$’s attributes deviate from the expected distribution $\mathcal{\boldsymbol{F}}_{w}$. Due to our black-box framework, the specific ${\mathcal{\boldsymbol{F}}}_{c}$ is unknown. Instead, we seek $x_{nw}$ that retains visual similarity but maximizes some collective spectral difference $\ell_{F}$ from $x_{w}$:
[…]
Maximizing the difference across all frequencies should lead to significant discrepancies in the values of ${\mathcal{\boldsymbol{F}}}_{c}$, forcing them to deviate from ${\mathcal{\boldsymbol{F}}}_{w}$. As semantic and non-semantic watermarks target different parts of the spectrum, we must craft a suitable attack strategy for each, identifying a proper loss function and optimization procedure.
5.1 Attacking Non-Semantic Watermarks
We now instantiate eq. (1) to defeat non-semantic watermarks, while §5.2 presents a method against semantic schemes. Following §4.1.1, non-semantic watermarks are in higher frequencies. Thus, we must disrupt this range without leaving visible cues. We present our high-frequency loss DFL and the visual loss $\ell_{d}$, explaining how each step updates the input to arrive at a watermark-free output.
Spectral Loss. We propose the Direct Fourier Loss (DFL):
$\forall\;x,y:\;\;\;\;\;DFL(x,\;y)=\|FT(x)-FT(y)\|_{1}$
where $FT$ is the ($2D$) Fourier transform. DFL calculates the band-wise spectral differences between $x$ and $y$, outputting the sum of their magnitudes. Optimizing DFL maximizes the amplitudes of the differences in the collective spectral representations between the sample being optimized and the watermarked image $x_{w}$, which leads to an output whose spectral magnitudes differ from $x_{w}$’s. Since the watermark is encoded in these amplitudes, it will be significantly disrupted, causing the detector to fail to recognize it. Yet, this is only guaranteed if the perturbation budget that controls the quality and similarity to the watermarked image is permissive enough for such significant distortions to occur. Luckily, this is true when disturbing high frequencies since the energy concentration in these bands, and accordingly, the watermark’s footprint (as it should be invisible) is naturally small. Thus, watermarks become volatile, as they can be interfered with via minute perturbations.
[…]
Norm: To further ensure no such peaks appear (especially for low resolutions), we use the $\|\|_{2}$ norm to place a geometric bound on the distance from the watermarked image, allowing us to slightly increase the budget used for LPIPS. Whenever LPIPS risks introducing such peaks, $\|\|_{2}$ grows, forcing the procedure to explore different directions. Yet, $\|\|_{2}$ alone is insufficient since geometric constraints cannot capture visual similarity and will either lead to sub-optimal solutions if used with extremely small thresholds or impact quality otherwise. This is especially true since we operate in a black-box framework without model feedback, potentially requiring larger perturbations, entailing that the perceptual loss should be accurate to preserve quality.
Final High-Frequency Destruction Procedure. (for non-semantic watermarks) Using the above components, we implement eq. (1). We only need to identify the rules for updating the sample to define our search space. Since we seek high-frequency disruptions that affect the rates at which individual pixels vary w.r.t. their neighbors, we should modify each pixel directly at each step. Thus, our algorithm accepts watermarked image $x_{w}$, LPIPS loss $\ell_{p}$, LPIPS threshold $t_{\ell_{p}}$, and $\|\|_{2}$ threshold $t_{\|\|_{2}}$, and finds modifier $\delta$ that maximizes DFL as follows:
[…]
$ReLU_{\ell}(x,\;y)$ denotes $max\{\ell(x,\;y)-t_{\ell},0\}$ for any loss $\ell$. Effectively, we instantiate $\ell_{d}$ from eq. (1) with two components: LPIPS loss $\ell_{p}$ and norm $\|\|_{2}$, each with its threshold, while DFL is the spectral loss. The vector $\delta$ captures the changes to $x_{w}$ to generate the non-watermarked $x_{nw}\!\!\!=\!\!x_{w}\!\!+\!\!\delta$. We minimize -DFL, maximizing DFL. The constants $c_{\ell_{p}}$ and $c_{\|\|_{2}}$ balance the spectral and perceptual losses. We enforce that each perceptual loss ($\ell_{d}$) component ($\ell_{p}$ and $\|\|_{2}$) remains below its threshold ($t_{\ell_{p}}\!$ or $t_{\|\|_{2}}$) through the term $ReLU_{\ell_{d}}(x_{w}\!\!+\delta,\;x_{w})$: When $\ell_{d}$ exceeds $t_{\ell_{d}}$, it evaluates to $\ell_{d}(x_{w}\!+\delta,\;x_{w})-t_{\ell_{d}}$, causing $\ell_{d}$’s gradients to produce a sample closer to $x_{w}$ in the next iteration. Otherwise, it returns 00, discarding $\ell_{d}$’s gradients and maximizing DFL unconditionally. $c_{\|\|_{2}}$ is a fixed large coefficient that strongly prevents the abovementioned peaks. $c_{\ell_{p}}$ is determined via a binary search: We start with a large $c_{\ell_{p}}$, assigning a larger weight to $\ell_{p}$’s gradients and run the optimization for several steps. If a solution is found, $c_{\ell_{p}}$ is decreased, increasing DFL’s dominance at the expense of larger visual deviations. Otherwise, $c_{\ell_{p}}$ is amplified. See Appendix A for details.
5.2 Attacking Semantic Watermarks
Semantic watermarks are in low frequencies, making them difficult to address. In addition to their large magnitudes (see §4.1.2), low frequencies correspond to many neighboring pixels varying gradually, indicating these watermarks are rooted in the image’s texture. Thus, removing them requires systemic and structural changes. Direct optimizations similar to §5.1 fail since they operate on the image and, at each step, individual pixels are updated in different directions. This targets higher frequencies, introducing significant changes yet only to a few pixels, failing to progress due to the perceptual constraints and incurring minimal changes at low frequencies. We could limit the procedure to optimize DFL over low-frequencies only. Yet, as significant changes are required in this range, without a methodical scheme, we will converge to a sub-optimal solution wherein a certain region receives substantial modifications, leading the perceptual loss to prevent further updates. This is because the process fails to account for inter-pixel links, and DFL does not spread the changes over the image.
(a) Basic: To calculate the output at $(i,\;j)$, the patch $PE^{i,j}_{\text{\tiny\mbox{$M\!{\times}\!N$}}}$ (where $M\!{\times}\!N$ is the kernel size) around $(i,\;j)$ is element-wise multiplied by the spectral kernel $\mathcal{K}$, summing over this product.
(b) Guided: The color kernel $\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}$ is first computed, assigning a modulation factor to each neighbor in the vicinity $PE^{i,j}_{\text{{\tiny\mbox{$M\!{\times}\!N$}}}}$ of $(i,\;j)$. These are element-wise multiplied by the spectral kernel $\mathcal{K}$ to limit the contribution from neighbors whose values differ significantly from $(i,\;j)$’s. This yields the effective kernel $\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma\!_{{}_{c}}}_{i,j}$ used to calculate the output at $(i,\;j)$ similar to basic filters.
Figure 3: Operation of known filters.
We develop a visually-aware attack inspired by image filtering. While filters apply systemic spectral disruptions, they fail to remove watermarks due to their content-agnostic nature, limiting the distortions they may introduce without leaving crucial visual cues. Our adversarial alternative retains quality by learning the filters’ weights to solve an optimization problem that maximizes low-frequency disruptions, resorting to a perceptual loss to preserve critical information. First, we motivate our design through guiding principles. We then provide technical details and present our optimizable filters. Finally, we introduce our attack.
5.2.1 Motivation
Basic Filtering: An image is convolved with a filter ($M{\times}N$ spectral kernel), moved s.t. its center is aligned with a single pixel $(i,\;j)$ at each step. The output at $(i,\;j)$ is calculated by multiplying the kernel by the pixel’s vicinity and summing over the product— see Fig. 3(a). The kernel’s values are non-negative and sum to $1$, making this output an averaged sum of the pixel and its neighbors. This performs noise reduction, decaying unwanted peaks. The kernel’s elements grow toward the center, assigning a larger weight to the original pixel to preserve the content.
Convolution in space translates to multiplication in the frequency domain: Given the spectra of the image and filter, it produces an image with a spectrum equivalent to their product. Thus, filters are good candidates for spectral manipulations. Yet, the outputs at any pixel are calculated using the same kernel. Hence, their potential to remove watermarks is limited [23, 35] due to their content-agnostic nature, making them only applicable with mild parameters as stronger blurring impacts quality.
Guided Filtering: These advanced filters tackle the information loss under stronger distortions, allowing more effective noise reduction while preserving crucial data. They are edge-preserving, minimizing the disruptions at these critical boundary points between different objects. Instead of using a uniform kernel, they involve an additional per-pixel color kernel that weakens the filter at edge points based on geometric principles: If some values in the pixel’s vicinity drastically differ from it, then it lies on an edge, as these neighbors belong to other components and the spectral kernel’s strength here must be diminished to preserve defining information. The color kernel computes a modulation factor for every neighbor in the filter’s vicinity around $(i,\;j)$ based on its value’s difference from $(i,\;j)$’s. The effective kernel for calculating $(i,\;j)$’s output is the Hadamard product of the two kernels— see Fig. 3(b). Guided filters fail to remove watermarks despite their stronger blurring (see §6.2.1) since they merely follow geometric heuristics that fail to fully capture human vision or target specific image characteristics, limiting the allowed distortion. In practice, less visually-critical areas can be changed drastically, regardless.
Adversarial Filtering (Our Attack): Based on the above, structure-aware filters can better maximize the spectral distortion with minimal impact on crucial regions. Similar to guided filters, we design a process for learning a set of per-pixel kernels. Yet, instead of relying on geometric heuristics, our method learns the kernels’ parameters s.t. they retain perceptual similarity to the original image from a human’s perspective, allowing for stronger filtering.
5.2.2 Formalizing Filters
Given image  Q×Rsuperscript $\mathcal{I}\!\!\in\!\!\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R$}}}}$, $PE^{i,j}_{\text{\tiny\mbox{$M\!{\times}\!N$}}}:\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R$}}}}$  M×Nabsentsuperscript $\!\!\longrightarrow\!\!\mathbb{R}^{\text{\tiny{ \mbox{$M\!{\times}\!N$}}}}$ denotes the local patch extractor that outputs the $M{\times}N$ vicinity around its $(i,\;j)^{th}$ pixel s.t. this pixel is at the center (i.e., $PE^{i,j}_{\text{\tiny\mbox{$M\!{\times}\!N$}}}(\mathcal{I})(\text{\small\mbox{$\left\lfloor\frac{M}{2}\right\rfloor$}},\;\text{\small\mbox{$\left\lfloor\frac{N}{2}\right\rfloor$}})=\mathcal{I}(i,\;j)$)— see Fig. 3(a). The global patch extractor $PE_{\text{\tiny\mbox{$M\!{\times}\!N$}}}:\mathbb{R}^{\text{\tiny{ \mbox{$Q{\times}R$}}}}\!\!\longrightarrow\!\!\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R{\times}M\!{\times}\!N$}}}}$ outputs a matrix that stores in cell $(i,\;j)$ the local extractor $PE^{i,j}_{\text{\tiny\mbox{$M\!{\times}\!N$}}}$. That is, $PE_{\text{\tiny\mbox{$M\!{\times}\!N$}}}(\mathcal{I})$’s $(i,\;j)^{th}$ cell contains the $M{\times}N$ patch around $(i,\;j)$ in $\mathcal{I}$. We formalize filters to present our solution.
[…]
Guided Filtering: These employ additional per-pixel color kernels to modulate distortions at critical points (Fig. 3(b)). Given kernel size $M{\times}N$, the color kernel for each $(i,\;j)$ is the matrix whose $(m,\;n)$ cell is $\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}(m,\;n)$ = $e^{\mbox{\small${-\frac{\|\mathcal{I}(i+m-\left\lfloor\frac{M}{2}\right\rfloor,\;j+n-\left\lfloor\frac{N}{2}\right\rfloor)-\mathcal{I}(i,\;j)\|_{1}}{2\sigma\!_{{}_{c}}^{2}}}$}}$. $\sigma\!_{{}_{c}}$ is a permissiveness controlling the modulation intensity. The $\|\|_{1}$ norm accounts for each pixel potentially having multiple color channels. For instance, RGB images associate each pixel with three values. The effective kernel for computing the output at $(i\;j)$ is then $\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma\!_{{}_{c}}}_{i,j}=\text{\small\mbox{$\frac{\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}\odot\mathcal{K}}{\underset{m,n}{\text{\large\mbox{$\sum$}}}\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}\odot\mathcal{K}}$}}$. At the center, $\mathcal{C}^{\mathcal{I},\sigma\!_{{}_{c}}}_{i,j}(\left\lfloor\frac{M}{2}\right\rfloor,\;\left\lfloor\frac{N}{2}\right\rfloor)=1$, indicating the contribution of the original $\mathcal{I}(i,\;j)$ is not decayed. For neighbors, large distances decrease their coefficients and limit their contributions. If we define $\mathcal{V}^{I,\mathcal{K},\sigma\!_{{}_{c}}}$ as the matrix whose $(i,\;j)^{th}$ cell contains $\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma\!_{{}_{c}}}_{i,j}$, the guided filter is given as $GF_{\mathcal{K},\sigma_{c}}(\mathcal{I})=\underset{m,n}{\text{\large\mbox{$\sum$}}}PE_{\text{\tiny\mbox{$M\!{\times}\!N$}}}(\mathcal{I})\odot\mathcal{V}^{\mathcal{I},\mathcal{K},\sigma\!_{{}_{c}}}$.
Optimizable Filters (Ours): Guided filters’ drawback is that they employ the same spectral kernel $\mathcal{K}$ for all pixels, changing only the color kernel computed via the same formula at all locations. Instead, we allow each pixel $(i,\;j)$ to have a unique spectral kernel ${\mathcal{K}\!\!\!_{\;\;i,j}}$, with weights that can differ significantly between locations based on feedback from a perceptual loss that determines their visual importance and allows maximizing the destruction at less critical regions. Our filters consist of a set $\mathcal{K}^{*}$ of local ${\mathcal{K}\!\!\!_{\;\;i,j}}$’s. We include color kernels to avoid exploring unwanted directions that violate vital geometric constraints. While these may be limiting, ignoring them causes unnatural distortions if the differences between neighboring pixels are extremely large, making them essential for remedying potential perceptual loss imperfections. When the distances are reasonably large but satisfy the perceptual loss with high confidence, it will assign larger weights that counteract the color kernel. Only when these distances are extremely large do they become prohibitive, which is ideal. Our optimizable filter is:
$\forall\mathcal{I}:\;OF_{\mathcal{K}^{*},\mathcal{R},\sigma\!_{{}_{c}}}(\mathcal{I})=\underset{m,n}{\text{\large\mbox{$\sum$}}}PE_{\text{\tiny\mbox{$M{\times}N$}}}(\mathcal{I})\odot\mathcal{V}^{\mathcal{R},\mathcal{K}^{*},\sigma\!_{{}_{c}}}$
$\mathcal{R}$ is a reference image for obtaining the color kernels, set to $\mathcal{R}\equiv x_{w}$. We now show how to learn $\mathcal{K}^{*}$’s weights.
5.2.3 Adversarial Filtering
[…]
where $\theta_{\mathcal{K}^{*}\!\!\!_{{}_{t}}}$ are the kernels’ weights and $\delta$ is a modifier that directly optimizes $x_{w}$, similar to eq. (2). We maximize the spectral loss by sequentially propagating the input through a network of filters. Recall that direct modifications are not effective for disrupting low frequencies in the absence of filters due to their inclination to cause high-frequency disturbances rather than systemic shifts in the values of many neighboring pixels in a gradual manner— see §5.2. Thanks to their convolutional nature, this no longer holds with our filters, making their outputs combinations of neighboring pixels. We still include $\delta$ that directly manipulates the input despite its inefficacy in the absence of filters. This is because the originally high-frequency disturbances caused by such modifications will now be distributed over neighbors of the receiving pixels due to our filters, translating them into useful low-frequency noise, causing further disruptions beyond the filters’ independent ability. We employ multiple ($T$) filters stacked after one another. The reason is that chaining several filters with various kernel sizes (geometries) targets different directions and slopes where the watermark could have been injected since watermarking is content-dependent and not all images can be changed similarly.
5.2.4 Spectral Loss
The former DFL is unsuitable for low frequencies (see §5.2) as it fails to force the changes to spread throughout the image. The collective energy at higher frequencies is much smaller in comparison, significantly reducing the total number of pixels that need to change to erase watermarks in these bands, which is achievable without violating the perceptual constraints. In contrast, sufficiently manipulating lower frequencies can rapidly change the structure, hindering convergence to an optimal solution if the distortion is not carefully distributed. This loss should also ensure our kernels perform the filtering operation, which DFL fails to enforce. Better alternatives are:
Filter Regularization Loss (FRL): It enforces that our kernels indeed operate as (smoothing) filters and is given as:
$&\forall x\in\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R$}}}}:\;\textit{\text{FRL}}(x)=-\sum_{t\in[\![T]\!]}\textit{\text{FRL}}_{{}_{{\boldsymbol{t}}}}(x)\\[-5.0pt] &\;\;\;\;\;\;\;\;\;\;\;s.t.\;\forall t:\;\;\textit{\text{FRL}}_{{}_{{\boldsymbol{t}}}}(x)=\|x-{{{{{\mu}}}}\!_{{}_{{}_{\frac{1}{2}}}}\!\!\!\!\!\!\!}^{{}^{{}^{\text{\tiny\mbox{$M_{t}\!{\times}\!N_{t}$}}}}}\!\!\!\!\!\!(x)\|_{2}\end{aligned}$
$M_{t}{\times}N_{t}$ is the size of our $t^{th}$ kernel ${\mathcal{K}^{*}\!\!\!_{{}_{t}}}$ and ${{{{{\mu}}}}\!_{{}_{{}_{\frac{1}{2}}}}\!\!\!\!\!\!\!}^{{}^{{}^{\text{\tiny\mbox{$M_{t}\!{\times}\!N_{t}$}}}}}\!\!\!\!\!\!$ computes the median of this vicinity around each pixel. Maximizing FRL minimizes $\textit{\text{FRL}}_{{}_{{\boldsymbol{t}}}}$ for every filter $t$, which is the sum of all pixels’ distances from the median in the filter’s environment around each pixel, bringing pixels closer to these local medians. This eliminates variations to the extent possible under the perceptual loss, leading to smoother outputs. As the watermark is distributed over the image, such patches will exhibit gradual “watermarking” variations that will now be lost. A major benefit of FRL is output consistency.
Mean Pool Loss (MPL): FRL is insufficient, ensuring only proximity of the pixels to the local medians but imposing no constraints on these medians. The means of the various patches in the watermarked image are highly influenced by the watermark that causes many pixels to change their values. Due to the perceptual loss that strives to retain information, the procedure may converge to a sub-optimal solution, skewing these medians toward the original means and only slightly affecting pixel intensities, retaining many shifts associated with the watermark. Thus, we require MPL:
$\forall x,y\in\mathbb{R}^{\text{\tiny{ \mbox{$Q\!{\times}\!R$}}}}:\;\;\textit{\text{MPL}}^{{}^{\text{\tiny\mbox{$L\!{\times}\!P$}}}}(x,\;y)=\|\overline{x}^{{}^{\text{\tiny\mbox{$L\!{\times}\!P$}}}}\!\!\!-\overline{y}^{{}^{\text{\tiny\mbox{$L\!{\times}\!P$}}}}\|_{1}$
$\overline{x}^{{}^{\text{\tiny\mbox{$L\!{\times}\!P$}}}}\!\!$ denotes mean pooling with an $L{\times}P$ window. MPL takes the means over each $L{\times}P$ patch in both inputs and calculates the $\ell_{1}$ difference. As most pixels do not encode watermark information, maximizing the means’ differences (w.r.t. the watermarked image) while demanding proximity to the medians (using FRL) primarily affects pixels located far (those that encode the watermark) from all others, bringing them closer to the rest and eliminating watermarking variations. A proper low-frequency loss must spread the changes over multiple regions (see §5.2.4). MPL does so by dividing the image into patches and demanding that each receive a portion of the change, implying the number of patches must be large. Yet, each must contain enough pixels to manipulate low frequencies. We choose ${\text{\small\mbox{{MPL}}}}^{{}^{\text{\tiny\mbox{$5\!{\times}\!5$}}}}\!\!$ as a compromise, balancing these requirements. Calculating MPL for all filter sizes similar to FRL is less desirable as when the sizes shift away from this sweet spot, one of the conditions is violated. Note that, without FRL, MPL is insufficient as it can lead to all pixels in each patch receiving a similar modification $\epsilon$, simply changing its brightness. This satisfies the perceptual constraints but leaves the spectrum unaffected as the pixel variation rates remain identical.
5.2.5 Final Low-Frequency Destruction Procedure
[…]
$x_{nw}=\;\underset{\{\theta_{\mathcal{K}^{*}\!\!\!_{{}_{t}}}\},\delta}{argmin}\;\left[\begin{array}[]{cc}c_{{{\text{\tiny\mbox{${DVL}$}}}}}\cdot ReLU_{{{\text{\tiny\mbox{${DVL}$}}}}}(x_{w},\;x_{w}^{{}^{OF_{{}_{\text{\tiny\mbox{$T$}}}},\delta}})+\lambda_{{\|\|_{2}}}\cdot\|x_{w}^{{}^{OF_{{}_{\text{\tiny\mbox{$T$}}}},\delta}}-x_{w}\|_{2}\\[5.0pt] -MPL^{{}^{\text{\tiny\mbox{$5{\times}5$}}}}(x_{w},\;x_{w}^{{}^{OF_{{}_{\text{\tiny\mbox{$T$}}}},\delta}})-\lambda_{{\text{\tiny\mbox{$FRL$}}}}\cdot FRL(x_{w}^{{}^{OF_{{}_{\text{\tiny\mbox{$T$}}}},\delta}})\end{array}\right]$
We maximize the loss $\ell_{F}\!=\!\!{\text{{MPL}}}^{{}^{\text{\tiny\mbox{$5{\times}5$}}}}\!\!\!\!+\lambda_{{\text{\tiny\mbox{$FRL$}}}}\cdot{\text{{FRL}}}$ by minimizing its negation similar to §5.1. $ReLU_{{{\text{\tiny\mbox{${DVL}$}}}}}$ ensures similarity, keeping the DVL distance from $x_{w}$ below $t_{{{\text{\tiny\mbox{${DVL}$}}}}}$. The resulting changes are gradual due to our filters, whose weights are learned to achieve this behavior thanks to our spectral loss, yielding optimal $\{\widehat{\theta}_{\mathcal{K}^{*}\!\!\!_{{}_{t}}}\},\widehat{\delta}$ s.t. the final solution is $x_{nw}\!\!=\!\!\text{\mbox{$x_{w}^{{}^{\widehat{OF}_{{}_{\text{\tiny\mbox{$T$}}}},\widehat{\delta}}}$}}\!\!\!\!$, where $\widehat{OF}_{{}_{\text{\tiny\mbox{$T$}}}}$ denotes the filters applied using $\{\widehat{\theta}_{\mathcal{K}^{*}\!\!\!_{{}_{t}}}\}$. As each kernel’s values should be non-negative and sum to $1$, we apply Softmax to the learned weights to obtain effective values in this range. $\lambda_{{\text{\tiny\mbox{$FRL$}}}}$ balances FRL and MPL and is calibrated s.t. both are optimized, while $\lambda_{{\|\|_{2}}}$ limits the impact of the $\|\|_{2}$ regularization. $c_{{\text{\tiny\mbox{$DVL$}}}}$ balances the spectral and visual losses and is determined via a binary search, as in §5.1. We found a single filter network architecture universally effective. Details are in Appendix A.
Discussion: Without our filters, the modifications at each step are less systemic, allowing different pixels to receive significant updates in opposing directions. Thus, we will converge to a sub-optimal solution since massive changes that a subset of pixels receive will substantially increase the visual loss, preventing further progress. While our low-frequency loss discourages this behavior and encourages uniform updates, it remains lacking, constantly countered by the visual loss that strives to retain similarity, limiting the number of pixels allowed to change their values, deeming such structural modifications less desirable compared to high-frequency changes that affect discrete pixels thereby remaining potentially unnoticeable. Our filters force prioritizing collective updates, drastically improving performance. Results establishing these benefits are in Appendix D.
While effective against semantic watermarks, this attack is unsuitable for non-semantic ones (see §6.2.2); The low-frequency loss and filters encourage low-magnitude, yet collective, changes and prevent significant updates to specific pixels (as opposed to §5.1). Yet, non-semantic watermarks are injected in high frequencies and thus rely on large variations between neighboring pixels. As this process can only slightly affect such peaks, the watermark is preserved.
6 Experiments
6.1 Setup
Schemes. We consider the SOTA watermarking schemes in Table I, covering all categories from §2, selected for appearing in top venues, with several remaining unbroken by a query-free black-box attacker. Among the generator-specific schemes, Yu1 [21], Yu2 [22], and PTW [23] watermark Generative Adversarial Networks (GANS) [53], while StableSignature [20] and TRW [18] operate on LDMs [46]. We are not aware of semantic schemes from top venues other than those we consider, which are the focus of previous work [35, 29, 31]. As most recent systems are generator-specific, we include five such schemes and two general-purpose ones. Compared to existing works (see §2.2), we span a wider variety of recent systems.
TABLE I: Studied SOTA watermarking schemes.
      General-Purpose Generator-Specific
Semantic     StegaStamp [24] TRW [18]
Non-Semantic     HiDDeN [19] Yu1 [21], Yu2 [22], PTW [23], StableSignature [20]
Data. Most generator-specific schemes require no input data. They were trained on various datasets: TRW and StableSignature on LAION-5B [54], Yu1 and Yu2 on CelebA [55], and PTW on FFHQ [56]. As TRW and StableSignature watermark text-to-image LDMs (StableDiffusion [2]), we use StableDiffusion’s prompts. For HiDDeN, we take samples from COCO [57] on which it was trained. StegaStamp watermarks $400{\times}400$ images. Thus, we downscale CelebA-HQ-1024 samples [^2] FFHQ-1024 samples yield similar results (excluded for brevity).. We evaluate each system on $100$ random images, similar to previous work [30].
Metrics. Following §3, we are concerned with two metrics:
1) Detection: Bit Accuracy. For binary watermarks, detection occurs if the ratio of bits matching the watermark in the sequence extracted from the image is above some threshold $t_{w}$. As in previous work [35, 32], we select $t_{w}$ that can reject the null hypothesis $H_{0}$ that $k$ or more matching bits were randomly extracted with $p<0.01$. This has probability $P(X\geq k|H_{0})=\sum_{i=k}^{n}\binom{n}{k}\frac{1}{2^{n}}$, for watermark length $n$. Thus, $t_{w}=\frac{k}{n}$ for which this probability is below $0.01$ is selected. We denote the portion of successfully extracted bits as the bit accuracy, which we compare against $t_{w}$. Watermark lengths are in Appendix B.
Inverse Distance. TRW’s watermarks are unbounded, and detection occurs if the $\ell_{1}$ distance (effectively, the Mean Absolute Error— MAE) of the extracted sequence from the watermark is below $t_{w}$ determined s.t. the ratio of falsely detected non-watermarked images (FPR) is low. We chose a higher FPR ($0.02$) than the original paper [18] ($0.01$), making the attacker’s task harder, obtaining $\textit{\text{MAE}}=71$. For consistency, we use this distance’s inverse: for the watermark to be detected, the MAE’s inverse must be $\geq t_{w}=\frac{1}{71}$.
Security Advantage. This binary indicator is the scheme’s ability to detect over $50\%$ of watermarked images. Most genuine content is not controversial, and detectors will not be invoked to determine its authenticity. They are only necessary when suspicious content surfaces. Absent additional knowledge, we assume a natural-looking controversial image has equal probabilities of being real or fake. Here, the best trivial detector that achieves the highest average accuracy on suspicious images would be a random classifier. An attacker will have a $50\%$ chance of fooling it with a fake image. For a scheme to have a Security Advantage, it must detect watermarked images at a higher rate. Otherwise, the resources invested in it will be in vain as it offers no additional protection against attackers compared to random systems. While such a scheme will perform better on non-watermarked images as it is trained to detect the existence of a watermark, this is irrelevant from a security aspect where only identifying the fake (watermarked) content matters.
2) Quality: We use two common  [35, 31, 32, 33, 23] human vision-inspired metrics: LPIPS [51] and Frèchet Inception Distance (FID) [58]. For fairness, we avoid LPIPS methods UnMarker explicitly optimizes (see §5.1), resorting to LPIPS-VGG [51]. FID compares the attack samples’ distribution to that of watermarked images, measuring dataset-wise similarity. We exclude geometric metrics (PSNR or norms) as we find them misleading. For instance, although visual metrics (Table III) and human inspection (Appendix E) verify the VAEAttack’s inferior quality, geometric methods deem it comparable with UnMarker.
Basic Manipulations. We evaluate Center Cropping ($10\%$), JPEG, Quantization, and Gaussian Blurring (filtering). Their parameters (see Appendix C) were chosen following previous work [23, 35] and to allow maximum distortion while retaining acceptable quality. We also evaluate Guided Blurring, showing how even these stronger filters cannot break robust watermarks, as argued in §5.2. To better explore these advanced filters, we even evaluate them with overly destructive parameters that do not represent viable attacks.
SOTA Attacks. We compare UnMarker against SOTA watermark removal attacks: The DiffusionAttack [35, 29, 31] and VAEAttack [35, 31]. To our knowledge, these regeneration methods are the only practical (query-free and black-box) and effective attacks, as adversarial attacks are disqualified for their unrealistic assumptions (see §2).
6.2 Evaluations
6.2.1 Watermark Utility
We ensure the chosen schemes’ utility, following the requirements from §4.1: Applicability, Stealthiness, and Robustness to Manipulations. While the latter demands withstanding basic manipulations, true robustness is contingent on resisting advanced attacks as well. We separate these cases for consistency with watermark papers that mainly consider basic manipulations. Advanced attacks are in §6.2.3. We focus on Applicability and Robustness as we use original implementations of SOTA systems whose Stealthiness has been shown by their authors.
Visual metrics are excluded here, as evaluating the naive manipulations’ output quality is not our objective, knowing that our basic settings are quality-preserving [35, 23]. We do not consider rotation manipulations, following Zhao et al.’s [35] arguments w.r.t. their impracticality as they can be easily spotted and corrected before invoking the detector. Still, their finding of some robust schemes vulnerable to rotations may seem at odds with our assertions that these watermarks are in the collective spectral amplitudes, which should leave them unaffected by rotations. Yet, schemes are generally not explicitly designed to construct watermarks in the spectral amplitudes. Rather, this behavior is adopted when trained for robustness (see §4). Failure to expose these systems to a wider set of demands (e.g., resistance to rotations) may result in them mistakenly associating additional indicators, such as a specific orientation, with the watermark. As rotations are impractical, this is not an issue.
TABLE II: Performance of watermarking schemes against each stage of UnMarker.
Note: Performance without attacks and under cropping is included for reference. The complete attack is in the last row (CHL). Numbers in parentheses are
the detection thresholds, while Detect is the portion of images meeting them. The best attacks are in bold font. All schemes lose their Security Advantage.
    Yu1 (61%) Yu2 (63%) HiDDeN (73%) PTW (70%) StableSig. (69%) StegaStamp (63%) TRW (0.0141)
[…]
None   98.42% 100% 99.71% 100% 99.16% 100% 97.96% 100% 98.97% 100% 99.95% 100% 0.022 100%
Cropping (10%)    74.56% 100% 53.71% 0% 86.84% 100% 93.9% 100% 97.86% 100% 93.96% 100% 0.0155 100%
UnMarker CH  59.04% 33% 53.06% 0% 58.34% 1% 63.11% 11% 53.08% 4% 94.87% 100% 0.0152 99%
CL  73.16% 100% 54.62% 0% 84.21% 98% 95.21% 100% 97.6% 100% 61.8% 43% 0.0141 55%
HL  73.11% 100% 99.23% 100% 72.28% 66% 61.53% 6% 54.59% 9% 64.64% 59% 0.016 98%
CHL  59.08% 31% 53.03% 0% 58% 0% 62.75% 11% 52.82% 4% 61.49% 43% 0.014 40%
A detailed analysis is in Appendix C alongside the results (Table IV). All schemes detect watermarked images without manipulations: For binary watermarks, all bit accuracies approach $100\%$. For TRW, the inverse distance is $0.022$, well over the $0.0141$ threshold. All systems are fairly resistant to manipulations. Semantic schemes remain almost unaffected. Among non-semantic schemes, Yu1 and StableSignature are the most resistant, retaining robustness against all realistic configurations. The latter shows some vulnerability to Guided Blurring with overly destructive parameters, but since this is not a practical attack, it remains reliable. PTW rejects most manipulations but is affected by reasonable Quantization ($10$), while the less advanced HiDDeN is vulnerable to JPEG and Blurring. Yu2 resists manipulations except cropping, indicating high dependency on spatial attributes, making it unsuitable for watermarking, as it violates the robustness requirement from §4.1.
The results demonstrate how all schemes except Yu2 are robust to spatial manipulations (e.g., cropping), indicating their watermarks are in a different channel, which we next verify is the spectral amplitudes as expected. This is evident for PTW and HiDDeN that exhibit vulnerability to Quantization and JPEG that affect high frequencies. The resistance of other schemes to these manipulations merely implies that their watermarks are harder to remove. Yet, advanced attacks can still break these systems, as shown below. Yu2 highly depends on information in the background, suggesting its watermarks are not constructed as robust, collective measurements but rather in volatile spatial locations.
6.2.2 UnMarker
We now evaluate UnMarker’s watermark removal capabilities. In §6.2.3, we compare it to SOTA attacks and assess its ability to preserve visual similarity.
Configuration. Our analysis above proves that a conventionally robust scheme (under spatial manipulations) must inject its watermarks in the frequency space. Schemes that are spatial in nature and rely less on spectral watermarking (e.g., Yu2) may hence evade UnMarker’s spectral optimizations as they are not designed to target these inherently fragile schemes. Yet, equipping UnMarker with a simple cropping layer suffices to defeat such weaker systems. Moreover, as spectral watermarks distribute their modifications throughout the image to minimize visual cues (see §4), cropping also can slightly “weaken” those by eliminating partial modifications residing in the background. This allows using a decreased distortion budget for the optimization stages.
Thus, UnMarker’s design is as follows: We apply mild cropping, followed by our high-frequency disruptions (§5.1). Finally, our filtering stage (§5.2) eliminates low-frequency traces. This yields a full attack that targets all carriers, erasing all watermarks. Attackers can adjust the cropping ratio or location to restrict the output to the actual content. For our experiments, $10\%$ center cropping sufficiently retains the content. The stages’ order does not affect the results.
UnMarker’s hyperparameters should be calibrated to maximize the spectral loss regardless of the system under attack, complying with our threat model. Hence, they are only affected by properties of the image, such as its size, which we found to be the only relevant factor. The hyperparameters are in Appendix A. The size-dependant parameters are merely the thresholds and coefficients. Meanwhile, the number of filters and their kernels are invariant, as by combining multiple filter sizes, a single architecture can defeat any semantic watermark (see §5.2).
Results. We report settings combining different stages of UnMarker, identifying crucial insights that establish our theory. These configurations are CH— the high-frequency destructive process preceded by cropping, CL— the low-frequency optimization (with cropping), HL— the two optimizations without cropping, and CHL, denoting the full UnMarker with all three stages. Results are in Table II. Following §6.1, we say that an attack defeats a scheme (alternatively, removes its watermark) if it causes it to lose its Security Advantage (i.e., detection below $50\%$).
Non-Semantic Schemes. Robust non-semantic schemes (w.r.t. known manipulations) cannot alter the content and thus must embed their watermarks in high frequencies. Hence, minimal disruptions to this part of the spectrum where the energy concentration is low suffices to remove them. Our results corroborate this as all such schemes break under CH that combines cropping and high-frequency disruptions. The detection rates of the conventionally robust algorithms (that resist cropping— all but Yu2) remain almost unaffected by CL, which primarily targets low frequencies where they do not operate. Yu2 is defeated by CH, but HL’s results prove this is not due to our optimization since the scheme is insensitive to manipulations in both parts of the spectrum, indicating that cropping only is responsible for this outcome. This comes as no surprise since Yu2 employs spatial watermarking that enables it to resist spectral optimizations but makes it vulnerable to cropping.
Schemes that process larger images (PTW and StableSignature) are hardly sensitive to cropping but vulnerable to high-frequency manipulations, shown by their robustness to CL and fragility against HL, bringing their detection to $6\%$ and $9\%$. Yu1 and HiDDeN that watermark smaller $128{\times}128$ images retain $100\%$ detection against HL. Yet, their bit accuracies drop sharply from $98.42\%$ and $99.16\%$ to $73.11\%$ and $72.28\%$. The cause is our high-frequency destruction since CL that combines cropping and low-frequency disruptions offers almost no advantages over cropping.
Thus, there is a connection between the image size, sensitivity to cropping, and vulnerability to high-frequency modifications, which is easy to explain: Spatial robustness and stealthiness require spreading the watermark across the image. Meanwhile, non-semantic watermarks are restricted to specific regions (i.e., edges) where they can exploit the high-frequency nature (see §4.1.1) to remain invisible. As the size shrinks, such areas become less prevalent in central parts, forcing the changes to be delegated to the background and making them more vulnerable to cropping. Accordingly, the portion of the change each region receives grows as well, making it difficult to eliminate these watermarks with minimal modifications, explaining why they retain bit accuracies above the threshold despite a significant decline without cropping. Regardless, mild cropping then becomes capable of eliminating sufficient footprints to further decrease these values below the thresholds, as seen for CH (CHL).
Larger images enable schemes to better reject cropping since this new-found resolution facilitates embedding the watermark in central areas, but the footprint in these regions must be smaller. Thus, minute disruptions throughout the image become more effective, explaining their inability to withstand high-frequency optimizations alone.
UnMarker (CHL) defeats all non-semantic schemes, with the most resistant being Yu1, which retains a detection rate of $31\%$. Yet, Yu1 still loses its Security Advantage (and so do all other schemes).
Semantic Schemes. StegaStamp and TRW resist high-frequency disruptions (CH), with their detection fixed at $100\%$ and the bit accuracies (inverse distances) retaining almost identical values to those observed for cropping alone. Yet, both are vulnerable to low-frequency disruptions, evident by their significant performance decline against CL: StegaStamp’s detection drops to $43\%$ while TRW’s becomes $55\%$. This cements our arguments regarding semantic watermarks being primarily in lower frequencies and the soundness of our filtering approach. Comparing CL to the final results (CHL), we observe identical detection for StegaStamp. TRW exhibits an additional decline to $40\%$, indicating a small portion of its watermark is in high frequencies.
StegaStamp’s watermarks are more visible (see §2.1), better engraving them in the main content, and increasing their robustness to spatial manipulations (cropping). Since TRW produces less visible watermarks despite targeting low frequencies where significant watermarking footprints are required, it behaves similar to non-semantic schemes for small images, forcing a considerable portion of these changes to populate the background. The results for cropping, CL, CH and CHL prove this; while cropping minimally affects StegaStamp (the bit accuracy only drops to $93.96\%$), TRW is more sensitive as its inverse distance decreases from $0.022$ to $0.155$ when the threshold is only $0.0141$. Although this is in the detection range, combining cropping with UnMarker’s low-frequency disruptions (CL, CHL) erases its watermarks (eliminates its Security Advantage).
Spectral optimizations (HL) drastically lower TRW’s inverse distance (to $0.016$), but do not erase its watermarks. This is because low-frequency distortions affect many pixels simultaneously. Under a strict budget, the number of regions where these rates can be altered is limited compared to high-frequencies. StegaStamp, whose watermarks are centralized, requires modifying fewer areas, making HL successful without cropping. Yet, we verified that low-frequency disruptions with a higher budget or number of steps break TRW without cropping, but this may introduce artifacts unless carefully done. Cropping is as effective but less demanding, confirming a flaw in TRW that targets non-critical regions.
TABLE III: Comparison with SOTA watermark removal attacks.
Note: Performance in the absence of attacks is included for reference. For each metric, bold entries correspond to the best attack against the relevant scheme.
Gray cells denote scores meeting our quality and removal criteria for LPIPS— 0.15 and Detect— $50\%$ (where the system loses its Security Advantage).
(a) Performance against non-semantic schemes.
    Yu1 (61%) Yu2 (63%) HiDDeN (73%) PTW (70%) StableSignature (69%)
[…]
None     NA NA 98.42% 100% NA NA 99.71% 100% NA NA 99.16% 100% NA NA 97.96% 100% NA NA 98.97% 100%
  UnMarker (CHL)     34.07 0.15 59.08% 31% 22.88 0.1 53.03% 0% 39.56 0.08 58% 0% 11.9 0.14 62.75% 11% 30.87 0.05 52.82% 4%
  Diffusion Attack no cropping     12.38 0.2 58.31% 28% 13.09 0.17 68.29% 87% 34.1 0.18 59.8% 3% 22.95 0.26 55.53% 7% 36.57 0.32 47.89% 0%
+cropping     14.02 0.18 54.24% 8% 11.98 0.17 51.26% 3% 36.97 0.19 56.61% 0% 19.98 0.25 62.01% 16% 41.88 0.31 47.78% 0%
  VAE Attack quality=3     36.2 0.15 67.19% 97% 19.67 0.13 73.89% 96% 65.16 0.22 60.4% 8% 26.6 0.26 73.63% 89% 72.21 0.37 48.25% 0%
quality=2     46.82 0.18 64.42% 81% 27.4 0.16 66.62% 78% 83.4 0.26 58.36% 1% 39.11 0.29 71.33% 79% 94.85 0.41 44.93% 0%
quality=1     58.37 0.22 60.63% 59% 38.9 0.19 61% 36% 104.92 0.32 56.26% 2% 54.93 0.33 66.74% 36% 128.88 0.44 43.87% 0%
(b) Performance against semantic schemes.
    StegaStamp (63%) TRW (0.0141)
[…]
None     NA NA 99.95% 100% NA NA 0.022 100%
  UnMarker (CHL)     34.69 0.08 61.49% 43% 60.73 0.1 0.014 40%
  Diffusion Attack no cropping     33.15 0.36 83.21% 100% 41.69 0.32 0.0155 99%
+cropping     30.83 0.33 77.84% 95% 41.16 0.31 0.0144 79%
  VAE Attack quality=3     54.89 0.42 97.9% 100% 71.91 0.38 0.0154 90%
quality=2     70.86 0.46 94.7% 100% 93.56 0.41 0.015 88%
quality=1     87.8 0.49 88.84% 100% 118.85 0.46 0.0145 67%
UnMarker defeats all semantic schemes due to irreconcilable trade-offs, as robustness to spatial manipulations incurs vulnerability to spectral disruptions and vice versa. Even when not bringing detection to $0\%$, it leaves a scheme useless, eliminating its Security Advantage.
Remarks: Any scheme either modifies the content (semantic) or does not (non-semantic). If attackers know the scheme’s class, they can apply only one of the stages (CH or CL), sufficing to erase the watermark, avoiding additional distortions. This information may be public or obtained by attacking multiple fake images (not the one intended for use in the actual attack), using the two stages independently, and later observing which images were not deleted (flagged). Some schemes may use principles from both classes. For instance, although TRW is semantic, a portion of its watermarks is in high frequencies, as discussed above. Here, applying both stages boosts the success rates. Yet, for all schemes (including less sophisticated ones, where only one stage is required), the full UnMarker (CHL) preserves quality even when both stages are applied (see §6.2.3).
6.2.3 Comparison with SOTA attacks
We compare UnMarker to regeneration attacks— the VAEAttack [35, 31] and DiffusionAttack [35, 29, 31]. Despite being certifiably guaranteed to remove non-semantic watermarks, they fail against semantic methods. This robustness creates a false sense of security around semantic watermarks [35, 31], which UnMarker disproves. Additionally, they guarantee the removal of non-semantic watermarks only if the attack introduces sufficient noise to dilute the watermark, followed by a denoising process to restore the image’s quality. Stronger watermarks require increased perturbations that prevent perfect reconstruction. Unnatural artifacts can be avoided by using advanced regeneration systems, explaining why the more sophisticated DMs outperform VAEs, as they can apply larger distortions while somewhat accurately retrieving the content. Yet, similarity to the original image may still be lost under such large budgets. UnMarker eliminates these hurdles, significantly decreasing the distortion while being effective even against the stronger semantic schemes.
As UnMarker uses cropping, we apply the DiffusionAttack with and without cropping for a fair comparison. The VAEAttack’s image quality was deemed inferior, disqualifying many of its results (see below), and thus a similar comparison was excluded. Results are in Table III, and samples produced by the three attacks are in Appendix E.
Quality: The range for LPIPS is $[0,1]$. Images with LPIPS distance at the $0.1$ level (i.e., below $0.2$) are visually similar [59]. Thus, we choose $0.15$ as a cut-off beyond which we say the attack begins to cause visible changes. This does not mean that slightly higher scores imply severe degradation since the changes may integrate well with the content. Yet, scores far beyond this threshold attest to poor output nature. Despite FID’s known ability to measure quality, it has two major shortcomings: first, it compares the datasets of watermarked images and those that have undergone the attack and cannot judge samples on an image-by-image basis. It may also exhibit inconsistencies with human vision, making it less ideal in certain instances [60]. Thus, while FID is important for assessing quality, we prioritize LPIPS.
We do not consider the VAEAttack effective. It allows choosing from eight quality options that control the added noise before regeneration. We apply it with the three lowest options, as we found increasing this factor beyond $3$ makes the attack almost always ineffective. Yet, all three options severely impact the quality, invalidating the attack. This can be seen in Table III, as both its FID and LPIPS are extremely worse (larger) than the other attacks whenever it degrades the detection rates. For larger images (over $256{\times}256$), even the best factor ($3$) is destructive, resulting in LPIPS scores at least 3.8×3.8$\times$3.8 × compared to UnMarker and well above the $0.15$ cut-off. Its FIDs are also inflated. While its FID against TRW ($71.91$) is close UnMarker’s ($60.73$), its LPIPS is much worse, indicating low quality. For smaller images, it retains similarity for Yu1, Yu2 and PTW with $quality\!\!=\!\!3$, but fails to remove their watermarks. HiDDeN’s watermarks are removed with this factor at an acceptable LPIPS ($0.22$). Still, the failure against other systems limits this attack. Lowering the factor impacts quality, invalidating any results for larger images (including PTW). For Yu1 and Yu2, as the attack fails to achieve remarkable removal for $quality\!\!=\!\!2$, the relevant case is $quality\!\!=\!\!1$. Both systems retain acceptable LPIPS scores ($0.22$ and $0.19$) but violate the $0.15$ threshold, and their FIDs are far worse than the other attacks, proving the VAEAttack fails to preserve quality (see Appendix E).
The DiffusionAttack is controlled via a parameter $t$. Following Saberi et al. [29], $t\!>\!0.2$ causes significant deviations from the watermarked input, invalidating the attack. We found that for smaller images ($128{\times}128$), this occurs at lower values. For these images, we set $t\!\!=\!\!0.05$, deeming it the turning point. A Lower $t$ results in less successful attacks. UnMarker achieves better LPIPS ($\leq\!\!0.15$), ensuring similarity. The differences are drastic for larger images (over $256{\times}256$)— $0.1$ at most for UnMarker compared to $\geq\!0.37$ for the DiffusionAttack. The higher LPIPS of DiffusionAttack, while not necessarily indicating significant deviations, entail that some nuances will appear, which may lead to its outputs being identified as fake— see Appendix E. The attacks’ FIDs alternate, occasionally outperforming one another. Yet, both attain far better FIDs than the VAEAttack.
UnMarker attains the lowest LPIPS, with its worst being $0.15$, ensuring resemblance to the watermarked images. The similarity between UnMarker’s FIDs and the DiffusionAttack’s, which is known to produce high-quality outputs [35], further establishes its capabilities.
Watermark Removal: Both SOTA attacks defeat non-semantic schemes, although the VAEAttack’s quality invalidates its results. UnMarker always outperforms this attack except against StableSignature for which UnMarker brings the detection to $4\%$ while the VAEAttack reduces it to $0\%$. Still, this difference is insignificant. The DiffusionAttack and UnMarker have similar results on average, but the DiffusionAttack (+cropping) significantly outperforms UnMarker on Yu1 ($8\%$ detection vs. $31\%$). Yet, as both eliminate its Security Advantage, specific numbers are irrelevant.
Our results validate regeneration’s known inability to remove semantic watermarks [35, 29, 31]. The VAEAttack moderately affects TRW, especially with quality $1$, bringing its detection to $67\%$, but, as explained above, we consider it invalid. The DiffusionAttack leaves both detection rates almost unchanged despite reducing StegaStamp’s bit accuracy and TRW’s inverse distance from $99.95\%$ and $0.022$ to $83.21\%$ and $0.0155$. While cropping decreases these numbers to $77.84\%$ and $0.0144$ and makes the attack more successful against TRW ($79\%$ detection), UnMarker drastically outperforms it, reducing this to $40\%$ and StegaStamp’s, against which DiffusionAttack is ineffective, to $43\%$.
Among the SOTA attacks, only the DiffusionAttack can remove watermarks without violating visual constraints. Yet, UnMarker is superior as it succeeds against semantic watermarks where the DiffusionAttack fails.
7 Discussion
Practicality. UnMarker is executed offline as it requires no detector interaction and takes up to $5$ minutes with a $40GB$ A100 GPU. Such machines are available for \$$\$32.77/hr$ on AWS. It requires no data, freeing it from the dependence on denoising pipelines (e.g., DMs). Regeneration attacks are affected by these systems’ fidelity and the commonalities between their training data and the watermarked images, as we showed for the VAEAttack. While the DiffusionAttack achieves better quality (but worse than UnMarker’s), its DM and all studied schemes were trained on public data. Providers possess exclusive data that differ from those used to train the attacker’s DM, making it likely to impact quality.
Mitigations. In the context of defensive watermarking, mitigations should allow recovery from removal attacks, enabling providers to continue using the scheme. Adversarial training (AT[61] is a popular mitigation in ML, wherein the model (detector) is trained on malicious samples from known attacks so that it learns to classify them correctly. Mitigations can also target the model’s inputs: the provider applies a pre-processing stage (e.g., purification [62]) to eliminate disruptions and only then invokes the model (detector). Mitigations could exist for adversarial attacks as they typically retain the sample’s true class but rely on model feedback to exploit its imperfections, yielding “volatile” perturbations in its uncertainty regions that cause misclassification. This strategy has also been the focus of previous work on watermark removal [32, 30, 29]. Yet, if the watermark is actually removed, recovery would be impossible. As we claim that by simply disrupting the carrier, UnMarker removes the watermark, mitigations should fail.
To test this, we consider a SOTA mitigation— diffusion purification [62]. Purification inspires the DiffusionAttack (see §6.2.3), operating similarly: It adds noise, diluting adversarial perturbations, and uses a DM for reconstructing a pure (attack-free) sample. As this applies as a certifiable attack on non-semantic watermarks, these schemes cannot use this mitigation. Yet, these watermarks are of little importance since these proofs for their removal undermine their practicality. As we aim to investigate the validity of defensive watermarking, we focus on the more reliable semantic schemes that meet the standards for utilizing purification as a defense. Still, UnMarker’s stages against both classes hinge on similar principles, making the findings applicable to defenses for non-semantic schemes if implemented.
We attack $100$ samples for each scheme using UnMarker (CHL). Then, we perform purification and invoke the detector. Given the resistance of semantic schemes to purification (the DiffusionAttack) and the effectiveness of this mitigation against adversarial examples, if UnMarker does not remove the watermark but overshadows it with volatile changes, purification would eliminate UnMarker and restore the watermark. Yet, the results verify UnMarker’s destructive nature that makes recovery impossible. For purification with $t\!\!=\!\!0.1$ (see §6.2.3), StegaStamp’s detection drops to $15\%$ and TRW’s to $16\%$. A larger $t$ leads to worse detection. Ironically, the defense degrades performance. While this is partly because diffusion may slightly decrease the bit accuracy (inverse distance), we apply it at half of its original strength (see §6.2.3), which allowed for $\sim\!\!100\%$ detection. The reason for this decline is UnMarker’s destruction, impairing the DM’s reconstruction ability. This raises the possibility of combining UnMarker with the DiffusionAttack as an even stronger attack, which we leave to future work due to the potential increased quality degradation. Evaluating AT is prohibitive due to time limitations, as it requires numerous adversarial examples, and a single UnMarker attack takes several minutes. Yet, AT is only successful if the attack sample retains the watermark, which is not the case for UnMarker, as proven. As purification adheres to similar principles, its ineffectiveness implies the same for AT.
Outlook. A natural question is whether watermarking can ever be made reliable. Particularly, can we devise schemes that do not conform to our spectral analysis s.t. UnMarker, or similar attacks, would fail? The answer is likely no.
In §4, we establish that a universal carrier that encodes the watermark is crucial for any scheme due to computational considerations. Given the nature of visual content, the assertion that this carrier must be in the spatial structure or spectrally constructed is also invariant. Then, we explain how spatial dependencies lead to fragile watermarks, as is demonstrated in §4.1. Thus, robust schemes must embed watermarks in the collective spectral amplitudes. Non-semantic schemes are unreliable and, despite their slightly demanding nature, regeneration attacks certifiably erase them, restricting strong watermarks to lower frequencies (semantic). While these are of larger magnitudes, they must remain minimally visible to retain quality and since obvious traces are removable even via naive editing. Existing semantic schemes push these demands to the extreme, leaving somewhat visible footprints (StegaStamp) or changing the content and limiting the space of watermarked images (TRW). Since UnMarker defeats them, future variants will have to further influence the content. Due to the considerable modifications the SOTA methods incur, such schemes will likely be impractical, reducing the quality and variety of watermarkable content and leaving visible traces. Combined with the inefficacy of mitigations, defensive watermarking becomes invalid, and future work should explore alternatives.
Limitations & Future Work. 1) We focus on practical (current) systems applicable to various images. Future schemes might be tailored for specific content (e.g., adding brand-specific glasses to face images), but such GenAI services will lack diversity. To meet users’ needs, providers will also offer less restrictive tools, and attackers can use them instead, rendering content-specific watermarking useless. To reduce visibility and resist naive attacks, they must also employ spectral manipulations (for reasons similar to §4.1.1’s), leaving UnMarker effective. Yet, without requiring universal applicability, the spectral footprint may target certain regions or objects only, not the collective spectrum. This necessitates augmenting UnMarker with content-based knowledge to identify optimal disruption regions, which could be studied should such schemes become prevalent. 2) UnMarker relies on accurate visual losses that preserve essential similarity cues while maximizing perturbations to attain the best removal rates, and future work may explore alternatives superior to ours (see §5.1). 3) While we consider visual content, our analysis translates to other domains (e.g., audio), but UnMarker requires domain-based adaptations that can be interesting to investigate. 4) As watermarks bear resemblance to adversarial examples (AEs), UnMarker could potentially defend against AEs. Still, as AEs do not require a universal carrier, they may target other channels (aside from collective spectra), mandating further improvements.
Ethical Considerations. Misusing GenAI has severe impacts on society. We aim not to help generate harmful content but raise awareness regarding these systems’ risks. Although we are not aware of providers that already released stable watermarking schemes, their adoption pace and attention from dominant entities (see §1) make assessments mandatory. We requested access to providers’ APIs for evaluation but never heard back. Thus, all studied systems are open-source.
8 Conclusion
We presented a novel attack against defensive watermarking via adversarial optimizations that disrupt the spectra of watermarked images and interfere with the watermarks’ footprints, erasing them without interacting with the scheme. UnMarker is effective against all systems, including semantic schemes that have been assumed robust to such query-free and black-box attacks. Our results show that defensive watermarking is not a viable countermeasure against deepfakes and that more robust defenses are needed.
Acknowledgments
We gratefully acknowledge the support of the Waterloo-Huawei Joint Innovation Laboratory for funding this research, and the Compute Canada Foundation (CCF) for their resources that made our experiments possible.
Appendix A UnMarker’s Hyperparameters
High-frequency Disruptions (Stage 1— §5.1): We use Alex [51] for small images ($128{\times}128$) and DVL [52] for larger sizes as the perceptual loss ($\ell_{p}$). As we maximize DFL, we set $t_{{\ell_{p}}}$ to the largest value at which similarity and quality are retained. This depends on $\ell_{p}$ and the image size. For small images optimized using Alex, this optimal point is $t_{{{\text{\tiny\mbox{${Alex}$}}}}}\!\!\!=\!\!5\!\cdot\!10^{-4}$. Since larger images require a different loss (DVL), similar thresholds do not indicate the same degree of resemblance. For $256{\times}256$ images, we find $t_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!4\!\cdot\!10^{-2}$ best achieves the above goals, while for larger images we have $t_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!\!10^{-4}$. Next, we select the initial $c_{{\ell_{p}}}$ that balances the spectral and visual losses. We want this starting value to be significant to assign a large weight to $\ell_{p}$ and ensure visual similarity is not violated, after which it can be decreased to allow the spectral loss to be more dominant while still ensuring this property. This initial setting only affects the time until it converges and must be empirically determined to speed up the process. We found an initial $c_{{{\text{\tiny\mbox{${Alex}$}}}}}\!\!=\!\!10^{6}$ optimal for small images. For DVL, this was set at $c_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!\!10^{6}$ when the size is above $256{\times}256$ as these images require the DVL loss to be extremely small ($t_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!\!10^{-4}$), while for $256{\times}256$ images, it can be far smaller ($c_{{{\text{\tiny\mbox{${DVL}$}}}}}\!\!=\!\!10^{-2}$).
The $\|\|_{2}$ component prevents unwanted peaks (see §5.1). Its threshold depends on the size and chosen LPIPS. While the LPIPS thresholds guarantee visual similarity to the extent that the images will look similar up to, potentially, such peaks, $\|\|_{2}$’s thresholds must eliminate these singular points when the LPIPS thresholds are met. We empirically selected these $\|\|_{2}$ thresholds: for $128{\times}128$, we have $t_{{\|\|_{2}}}=10^{-4}$, while $256{\times}256$ images allow $t_{{\|\|_{2}}}=3\cdot 10^{-5}$, and larger images operate at $t_{{\|\|_{2}}}=10^{-4}$. We report the normalized thresholds (divided by the image size). In practice, they increase with the size since larger images accommodate stronger perturbations without visible cues. $c_{{\|\|_{2}}}$ multiplies this $\|\|_{2}$ when its threshold is violated to strongly steer the procedure away from these boundaries. This is not determined through a binary search, as we want it to have the same effect constantly, without undoing the previous iterations’ effects. By tracking the losses, we arrive at $c_{{\|\|_{2}}}=0.6$.
TABLE IV: Performance of watermarking schemes against basic manipulations.
Note: The first row shows the results in the absence of any manipulations. Numbers in parentheses are the detection thresholds, while Detect is the portion of
the images meeting them. For attacks with multiple configurations, lower rows correspond to stronger distortions. * denotes overly-destructive configurations.
    Yu1 (61%) Yu2 (63%) HiDDeN (73%) PTW (70%) StableSig. (69%) StegaStamp (63%) TRW (0.0141)
[…]
None   98.42% 100% 99.71% 100% 99.16% 100% 97.96% 100% 98.97% 100% 99.95% 100% 0.022 100%
  Cropping ratio = 10%  74.56% 100% 53.71% 0% 86.84% 100% 93.9% 100% 97.86% 100% 93.96% 100% 0.0155 100%
  Blur std. = 0.5  98.13% 100% 99.55% 100% 96.45% 100% 96.93% 100% 98.52% 100% 99.95% 100% 0.0221 100%
std. = 1.0  94.6% 100% 99.73% 100% 62.15% 5.00% 82.25% 100% 86.65% 99% 99.88% 100% 0.0207 100%
  Guided Blur color std. = 0.1  87.91% 100% 99.32% 100% 75.6% 76% 94.41% 100% 96.38% 100% 99.91% 100% 0.0212 100%
color std. = $1.0^{*}$  85.32% 100% 99.43% 100% 56.12% 0% 78.66% 98% 71.25% 69% 99.93% 100% 0.02 100%
color std. = $5.0^{*}$  84.9% 100% 99.62% 100% 53.81% 0% 76.42% 92% 66.54% 50% 99.93% 100% 0.0199 100%
  Quantization factor = 8  85.86% 100% 99.16% 100% 89.34% 98% 74.92% 93% 98.83% 100% 99.83% 100% 0.0207 100%
factor = 10  82.51% 100% 98.7% 100% 86.51% 98% 68.34% 46% 99.04% 100% 99.8% 100% 0.0201 100%
  JPEG quality = 110  89.58% 100% 99.55% 100% 70.73% 43% 87.66% 100% 97.68% 99% 99.94% 100% 0.0215 100%
quality = 80  76.38% 100% 93.15% 100% 66.48% 16% 72.17% 76% 90.66% 100% 99.89% 100% 0.202 100%
By monitoring the spectral loss, we can determine when it stops increasing for each value of $c_{{\ell_{p}}}$ and the number of steps after which updating $c_{{\ell_{p}}}$ does not result in improvements. This yields $5000$ iterations for images smaller or larger than $256{\times}256$, and $2000$ otherwise. The number of binary steps is $2$ for sizes $\leq$ $256{\times}256$ and $4$ otherwise. The optimizer is Adam with learning rate $2\cdot 10^{-4}$, reduced by $0.5$ when no improvement (above $10^{-4}$) occurs for $10$ iterations. We clip gradients at $5\cdot 10^{-3}$ for DVL and $5\cdot 10^{-2}$ for Alex to ensure they do not dominate the process.
Low-frequency Disruptions (Stage 2— §5.2): Each filter is characterised by the dimensionality $M_{t}{\times}N_{t}$ of ${\mathcal{K}^{*}\!\!\!_{{}_{t}}}$ and permissiveness $\!\!{\sigma\!_{c_{{}_{t}}}}$. We find ${\sigma\!_{c_{{}_{t}}}}\!\!=\!\!5\cdot 10^{-2}$ optimal. When choosing the kernel sizes and number of filters $T$, resources may be prohibitive as we have a kernel per pixel, making a filter scale as the number of pixels multiplied by its kernel size. We found $T\!\!\!=\!\!9$ with different sizes defeats all studied systems. This architecture was chosen (semi-randomly) to target various potential geometries of the watermark under our computational limitations ($40$GB A100 GPU) but can be extended with better hardware. Yet, our findings indicate that this may not be necessary. The ordered sizes are $(21,5),\!$ $(5,5),(17,33),(7,7),(47,5),(33,17),(17,17),(5,5)\!\!\!$ and $(3,3)$. Our spectral loss consists of MPL and FRL. While we maximize both, FRL is a supporting regularization term— see §5.2.4. Hence, we deem a sample optimal if it maximizes MPL. Yet, $\lambda_{{\text{\tiny\mbox{$FRL$}}}}$ must account for FRL, ensuring the outputs meet the conditions it enforces. $\lambda_{{\text{\tiny\mbox{$FRL$}}}}\!\!=\!\!5$ was found optimal, but it can be adjusted if the two components do not meet expectations (see §5.2.4). We found $\lambda_{{\|\|_{2}}}\!\!=\!\!2.5\cdot 10^{-4}$ ideal but it can be calibrated similarly.
We universally use DVL (see §5.1). As low-frequency perturbations do not introduce sporadic peaks, the remaining parameters can be selected uniformly with little regard to the image size. The exception is the threshold $t_{{{\text{\tiny\mbox{${DVL}$}}}}}$, as smaller images ($\leq 256{\times}256$) are more sensitive to structural changes (unlike high-frequency perturbations that these images’ low resolution masks), exhibiting visible differences for $t_{{{\text{\tiny\mbox{${DVL}$}}}}}>10^{-3}$, while larger images can withstand 3×3$\times$3 × this change. With an initial $c_{{{\text{\tiny\mbox{${DVL}$}}}}}=10^{6}$, a single step running for $500$ iterations is sufficient. We use Adam with a $10^{-2}$ learning rate for the filters and $2\cdot 10^{-3}$ for $\delta$, decayed by $0.99$ if MPL fails improve by $5\cdot 10^{-2}$ for $10$ iterations. As DVL’s impact is weakened due to this procedure’s structural nature, we allow its gradients to be more dominant, clipped at $1.0$.
Filters operate on all image channels identically. Hence, as the reason behind including the modifier $\delta$ is the filters’ ability to smoothen its otherwise high-frequency changes, $\delta$ must also affect all channels similarly. Thus, we use a gray-scale modifier $\delta$ that applies the same changes to all channels. While manipulating spectral amplitudes impacts the texture and results in less critical changes, phase modifications are more visible (see §4.1.1). Hence, we should ideally restrict our filters to manipulate magnitudes. The frequency response of a filter does not change the phase iff it is even-symmetric (identical to the reflections of each of its quadrants across both axes). While this holds for basic (convolutional) filters, ours will converge to similar kernels for neighboring pixels that are of equal visual importance and can be altered similarly, indicating that each region will behave as if it were “convolved with a basic kernel”. Thus, our filters should possess this symmetry. Accordingly, instead of optimizing all weights, we focus only on the first quadrant of each kernel and reflect it across both axes.
Appendix B Watermarking Schemes’ Configuration
The (author-reported) watermark lengths are in Table V.
TABLE V: Watermark lengths.
Note: TRW’s watermarks are not binary; thus, this metric is not applicable.
 Scheme Length (# bits)
 Yu1 128
Yu2 100
HiDDeN 30
PTW 40
StableSignature 48
StegaStamp 100
TRW NA
Appendix C Baseline Results Against Image Manipulations
Figure 4: Outputs of the three removal attacks. The VAEAttack over-smooths images, resulting in them losing crucial information and occasionally looking cartoonish. Compared to the DiffusionAttack, UnMarker’s outputs better resemble the watermarked images. The DiffusionAttack also introduces semantic incoherences such as additional teeth in the case of PTW and omits identifying details such as freckles on the forehead for StegaStamp. PTW’s watermarked image itself is of inferior quality with an unnatural patch. The DiffusionAttack retains this artifact while UnMarker eliminates it, enhancing the image.
The kernel size for Gaussian (Guided) Blurring is $5{\times}5$. We exploit Guided Blurring’s edge-preserving properties to increase the std. (to $5.0$). To control the modulation at edges, these accept an additional color kernel’s std. (see §5.2), for which we have three values: 0.1, 1.00.11.00.1,\;1.00.1 , 1.0, and $5.0$. Yet, $1.0$ and $5.0$ are overly destructive and do not preserve quality, chosen only to explore robustness. All parameters are in Table IV.
Results: Most systems are robust to all practical manipulations. Despite resisting all other modifications, Yu2 breaks completely under cropping. Hence, unlike other schemes, it highly depends on spatial attributes [^3] Yu et al. [22] only evaluate a single ProGAN [55] architecture under manipulations, whereas we chose their StyleGAN2 [63] model.. Yu1, HiDDeN and TRW exhibit a considerable drop in the bit accuracies (inverse distance) after cropping, although their overall detection is not affected. This is not surprising, as the watermarking spectral modifications must be spread throughout the image to minimize visual cues, implying that a portion may be allocated to the background. Yet, they retain robustness by restricting most of these changes to central regions.
Yu1, StegaStamp and TRW resist all manipulations, while StableSignature can withstand realistic, non-obtrusive modifications, making it also robust for practical purposes. PTW is slightly less powerful, losing some of its efficacy under even non-destructive operations, such as Quantization with a factor of $10$ and JPEG with quality $80$ that decrease its performance to $46\%$ and $76\%$, respectively. HiDDeN exhibits similar behavior, resisting all realistic manipulations aside from JPEG that brings its detection to $43\%$ even when used with a reasonably good quality of $110$ or Blurring with $std.=1.0$ that lowers its detection score to $5.0\%$.
It is expected that some non-semantic systems (e.g., PTW, HiDDeN and StableSignature) begin to break under spectral manipulations such as JPEG and Blurring. As explained in §4.1, constructing spectral non-semantic watermarks should be done in an input-specific manner that is also persistent, which is a challenging task that will be restricted to volatile watermarks unless the scheme is sufficiently advanced. This explains why the more sophisticated StableSignature and PTW outperform HiDDeN. Yet, allowing for severe manipulations via stronger (Guided) Blurring breaks StableSignature as well, although this is only included to demonstrate this phenomenon that attests to the spectral nature of these watermarks and is not a valid attack (see §6.1). Quantization can cause sharp pixel shifts (high-frequency noise), explaining PTW’s vulnerability.
With few exceptions, which we attribute to a slight mismatch in the experimental settings, our findings align with those attained by the authors of the schemes that evaluate the same attacks with similar parameters.
Appendix D Effects of Adversarial Filtering
TABLE VI: UnMarker’s performance against semantic schemes with and without optimizable filters.
    StegaStamp (63%) TRW (0.0141)
[…]
UnMarker with filters  34.69 0.08 61.49% 43% 60.73 0.1 0.014 40%
w/o filters  30.45 0.04 93.13% 100% 49.59 0.04 0.0147 86%
The reader may wonder whether the low-frequency loss alone can defeat semantic schemes, eliminating the need for our filters. Based on our arguments explaining how direct optimizations are of limited applicability against low-frequency watermarks due to their non-systemic nature (see §5.2), this is not the case. For validation, we compare UnMarker against semantic schemes with and without our filters. In both experiments, the loss is the same from §5.2.4. The results in Table VI prove the filters necessary. While TRW experiences a slight degradation without filters, including them is far more effective, bringing its detection to $40\%$ (compared to $86\%$ otherwise). StegaStamp is unaffected when filters are excluded, retaining $100\%$ detection. With filters, this drops to $43\%$. We also verified the filter-free attack’s ineffectiveness persists even if the perturbation slightly increases. Although similarity is slightly worsened by our filters as they encourage structural changes, the scores show that, even with filtering, UnMarker retains quality.
Appendix E Outputs of UnMarker vs. The SOTA
Fig. 4 displays examples of watermarked images and the outputs obtained by subjecting them to advanced attacks.
Appendix F Meta-Review
The following meta-review was prepared by the program committee for the 2025 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
F.1 Summary
This paper introduces UnMarker, a universal attack method designed to compromise defensive image watermarking mechanisms.
F.2 Scientific Contributions
  • Identifies an Impactful Vulnerability
  • Provides a Valuable Step Forward in an Established Field
  • Independent Confirmation of Important Results with Limited Prior Research
F.3 Reasons for Acceptance
  1. 1.
    UnMarker demonstrates effectiveness against multiple watermarking schemes.
  2. 2.
    The paper provides valuable insights into the vulnerabilities of current watermarking methods by leveraging spectral amplitude analysis.